# Direct Adaptive Control of Grid-Connected Power Converters via Output-Feedback Data-Enabled Policy Optimization

### v1

@@ Abstract
Power electronic converters are gradually becoming the main components of modern power systems due to the increasing integration of renewable energy sources. However, power converters may become unstable when interacting with the complex and time-varying power grid. To deal with this problem, an adaptive control design scheme for power converters is preferable, which can capture the closed-loop dynamical interaction between the converter and the grid via online data. In this paper, we propose an adaptive data-driven control method, called data-enabled policy optimization (DeePO), to stabilize power converters by using only online input-output data. Our contributions are threefold. First, we propose a covariance parameterization of partially observed linear systems with input-output data. Second, we develop a DeePO algorithm, which updates the parameterized policy with data-based gradient descent to achieve computationally efficient adaptive control. Third, we use high-fidelity simulations to verify that DeePO can effectively stabilize grid-connected power converters and quickly adapt to the changes in the power grid.
@@ I Introduction
Recently, there has been a renewed interest in direct data-driven control, which bypasses the system identification (SysID) step and learns the controller directly from a batch of persistently exciting data [10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. This approach is end-to-end, easy to implement, and has seen many successful applications [20, 21, 22]. Following this line, our previous work proposes a direct data-driven linear quadratic regulator (LQR) method [23, 24]. It is adaptive in the sense that the control performance is improved in real time by using online closed-loop data. We call this method Data-enabled Policy Optimization (DeePO), where the policy is parameterized with sample covariance of input-state data and updated using gradient methods. DeePO has a recursive algorithmic implementation and hence is computationally efficient. From a theoretical perspective, it meets provable stability and convergence guarantees for linear time-invariant systems. All these key features render DeePO an ideal data-driven control method for power converters.
In this paper, we propose a DeePO method to mitigate oscillations in power converter systems. Since the power converter system is only partially observed, i.e., the full state is unmeasurable, we first extend the covariance parameterization of the system with input-state data in [23, 24] to that with input-output data. This is achieved by regarding a finite length of past input-output trajectory as the state and using the corresponding covariance to parameterize the closed-loop system. Then, we propose the DeePO algorithm for direct adaptive control of partially observed systems. Finally, we apply the DeePO algorithm to stabilize power converter and direct-drive wind generator systems, both of which are unknown, state-unmeasurable, and encounter sudden change of dynamics due to grid changes. Simulation results show that DeePO enables efficient online adaptation and effectively prevents instabilities in grid-connected power converters. Compared with our previous works applying data-enabled predictive control (DeePC) to power converters [21, 25], DeePO offers a significantly reduced online computational burden, making it more suitable for scenarios where the processor cannot solve a quadratic program in real time.
The rest of the paper is organized as follows. Section II formulates the stabilization problem of power converter systems. Section III introduces the adaptive LQR control and the DeePO method. Section IV extends the DeePO algorithm to input-output systems. Section V performs simulations on the converter systems. Conclusion is made in Section VI.
@@ II-A Power converter systems
where $x_{t}\in\mathbb{R}^{n}$ is the state variable (could be unmeasurable, e.g., variables in the power grid side), $u_{t}\in\mathbb{R}^{m}$ is the control input (e.g., additional signals added to the current references; see Fig. 1), $y_{t}\in\mathbb{R}^{d}$ is the output (e.g., active and reactive power in our setting), $w_{t}$ is the process noise, and $v_{t}$ is the measurement noise. The unknown system $(A,B,C)$ is controllable and observable, but it may be subject to sudden impedance changes due to external grid changes, e.g., changes of $Z_{\rm grid}$ in Fig. 1. Such events may lead to poorly damped or even unstable oscillations.
@@ II-B Challenges in stabilization of power converter systems (DC source + DC-AC converter + AC power grid)
  • unknown model: Power converter systems often exhibit high-order characteristics, rendering it difficult to establish exact dynamical models. Specifically, the exact dynamics of the DC source and the converter are usually hard to obtain (and therefore unknown) due to proprietary manufacturer models and their complexity, e.g., the complex mechanical behaviors of wind turbines. Moreover, the power grid is also unknown when designing a controller for the converter, as the grid is large-scale and time-varying (therefore hard to model).
@@ II-C Our approach
In this paper, we propose a direct data-driven method to solve the stabilization problem of power converter systems without SysID. Our approach is based on data-enabled policy optimization (DeePO), an adaptive linear quadratic control method introduced in our recent work [23, 24]. It is data-driven and does not involve any explicit SysID. Moreover, DeePO uses online closed-loop data to adaptively update the control policy, making it an ideal method to deal with time-varying dynamics.
@@
In the sequel, we first recapitulate the DeePO method with input-state data in [23, 24]. Then, we extend DeePO to the input-output setting. Finally, we discuss its implementation on power converter systems and perform simulations to validate the effectiveness of DeePO.
@@
III Direct adaptive LQR control via data-enabled policy optimization
In this section, we first introduce the linear quadratic regulator (LQR). Then, we recapitulate data-enabled policy optimization for adaptive learning of the LQR [23, 24].
III-A The linear quadratic regulator
Consider a linear time-invariant system with fully measurable state
@@ III-A The linear quadratic regulator
Here, $h_{t}$ is the performance signal of interest, $(A,B)$ is controllable, and the weighting matrices $(Q,R)$ are positive definite.
The LQR problem is phrased as finding a state-feedback gain $K\in\mathbb{R}^{m\times n}$ that minimizes the $\mathcal{H}_{2}$-norm of the transfer function $\mathscr{T}(K):w\rightarrow h$ of the closed-loop system
$\begin{bmatrix}x_{t+1}\\ h_{t}\end{bmatrix}=\begin{bmatrix}A+BK&I_{n}\\ \hline\cr\begin{bmatrix}Q^{1/2}\\ R^{1/2}K\end{bmatrix}&0\end{bmatrix}\begin{bmatrix}x_{t}\\ w_{t}\end{bmatrix}.$
When $A+BK$ is stable, it holds that [26]
$\|\mathscr{T}(K)\|_{2}^{2}=\text{Tr}((Q+K^{\top}RK)\Sigma_{K})=:C(K),$ (3)
@@
$\Sigma_{K}=I_{n}+(A+BK)\Sigma_{K}(A+BK)^{\top}.$ (4)
We refer to $C(K)$ as the LQR cost and to (3)-(4) as a policy parameterization of the LQR.
It is well-known that the optimal LQR gain $K^{*}$ is unique and can be found by, e.g., solving an algebraic Riccati equation with $(A,B)$ [26]. When $(A,B)$ is unknown, data-driven methods learn the LQR gain from input-state data. In the sequel, we recall the DeePO approach proposed in our previous work [23, 24].
III-B DeePO for direct adaptive LQR control
Policy optimization (PO) refers to a class of direct design methods, where the policy is parameterized and recursively updated using gradient methods [27]. In particular, DeePO uses the sample covariance of input-state data to parameterize the policy [24]. Consider the $t$-long time series of states, inputs, noises, and successor states
$X_{0,t}$ $:=\begin{bmatrix}x_{0}&x_{1}&\dots&x_{t-1}\end{bmatrix}\in\mathbb{R}^{n\times t},$
@@ III-B DeePO for direct adaptive LQR control
$W_{0,t}$ $:=\begin{bmatrix}w_{0}&w_{1}&\dots&w_{t-1}\end{bmatrix}\in\mathbb{R}^{n\times t},$
$X_{1,t}$ $:=\begin{bmatrix}x_{1}&x_{2}&\dots&x_{t}\end{bmatrix}\in\mathbb{R}^{n\times t},$
@@
$X_{1,t}=AX_{0,t}+BU_{0,t}+W_{0,t}.$ (5)
Assuming that the data is persistently exciting (PE) [28], i.e., the block matrix of input and state data $D_{0,t}:=[U_{0,t}^{\top},X_{0,t}^{\top}]^{\top}$ has full row rank
$\text{rank}(D_{0,t})=m+n,$ (6)
the covariance parameterization[24] of $K$ is given by
$\begin{bmatrix}K\\ I_{n}\end{bmatrix}=\frac{1}{t}D_{0,t}D_{0,t}^{\top}V=:\Phi_{t}V,$ (7)
where $\Phi_{t}\succ 0$ denotes the sample covariance of input-state data, and $V\in\mathbb{R}^{(n+m)\times n}$ is the parameterized policy.
With the covariance parameterization (7), the LQR problem (3)-(4) can be expressed by raw data matrices $(X_{0,t},U_{0,t},X_{1,t})$ and the optimization matrix $V$. For brevity, let $\overline{X}_{0,t}=X_{0,t}D_{0,t}^{\top}/t$ and $\overline{U}_{0,t}=U_{0,t}D_{0,t}^{\top}/t$ be a partition of $\Phi_{t}$, and let $\overline{W}_{0,t}=W_{0,t}D_{0,t}^{\top}/t$ be the noise-state-input covariance, and finally define the covariance with respect to the successor state $\overline{X}_{1,t}=X_{1,t}D_{0,t}^{\top}/t$, Note that the dimension of the covariance matrices does not depend on $t$. Then, the closed-loop matrix can be written as
@@
Following the certainty-equivalence principle [14], we disregard the unmeasurable $\overline{W}_{0,t}$ and use $\overline{X}_{1,t}V$ as the closed-loop matrix. After substituting $A+BK$ with $\overline{X}_{1,t}V$ in (3)-(4) and leveraging (7), the LQR problem becomes
$\mathop{\text{minimize}}\limits_{V}~{}J_{t}(V):=\text{Tr}\left((Q+V^{\top}\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}V)\Sigma_{t}(V)\right),$ (8)
subject $\text{subject to}~{}~{}\overline{X}_{0,t}V=I_{n},$
where $\Sigma_{t}(V)=I_{n}+\overline{X}_{1,t}V\Sigma_{t}(V)V^{\top}\overline{X}_{1,t}^{\top}$ is a covariance parameterization of (4), and the gain matrix can be recovered by $K=\overline{U}_{0,t}V$. We refer to (8) as the direct data-driven LQR problem, which does not involve any explicit SysID.
The DeePO algorithm uses online gradient descent of (8) to recursively update $V$. The details are presented in Algorithm 1, which alternates between control (line 2) and policy update (line 3-5). The initial stabilizing policy can be either user-specified or computed by solving (8) with offline data. At time $t$, we apply the linear state feedback policy $u_{t}=K_{t}x_{t}+e_{t}$ for control, where $e_{t}$ is a probing noise used to ensure the PE rank condition (6). We use online projected gradient descent in (10) to update the parameterized policy, where the projection $\Pi_{\overline{X}_{0,t+1}}:=I_{n+m}-\overline{X}_{0,t+1}^{\dagger}\overline{X}_{0,t+1}$ onto the nullspace of $\overline{X}_{0,t+1}$ is to ensure the subspace constraint in (8). The gradient can be computed by the following lemma. Define the feasible set of (8) (i.e., the set with stable closed-loop matrices) as $\mathcal{S}_{t}:=\{V\mid\overline{X}_{0,t}V=I_{n},\rho(\overline{X}_{1,t}V)<1\}$.
Lemma 1 ([24])

### v2

@@ Abstract
Power electronic converters are becoming the main components of modern power systems due to the increasing integration of renewable energy sources. However, power converters may become unstable when interacting with the complex and time-varying power grid. In this paper, we propose an adaptive data-driven control method to stabilize power converters by using only online input-output data. Our contributions are threefold. First, we reformulate the output-feedback control problem as a state-feedback linear quadratic regulator (LQR) problem with a controllable non-minimal state, which can be constructed from past input-output signals. Second, we propose a data-enabled policy optimization (DeePO) method for this non-minimal realization to achieve efficient output-feedback adaptive control. Third, we use high-fidelity simulations to verify that the output-feedback DeePO can effectively stabilize grid-connected power converters and quickly adapt to the changes in the power grid.
@@ I Introduction
Recently, there has been a renewed interest in direct data-driven control, which bypasses the system identification (SysID) step and learns the controller directly from a batch of persistently exciting data [10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. This approach is end-to-end, easy to implement, and has seen many successful applications [20, 21, 22]. Following this line, our previous work proposes a direct data-driven linear quadratic regulator (LQR) method [23, 24]. It is adaptive in the sense that the control performance is improved in real time by using online closed-loop data. We call this method Data-enabled Policy Optimization (DeePO), where the policy is parameterized with sample covariance of input-state data and updated using gradient methods. DeePO is computationally efficient, meets provable stability and convergence guarantees for linear time-invariant systems, and has successful real-world applications [25]. However, the DeePO method [24] works only on the state-feedback control problem, which is not the case for the power converter system.
In this paper, we propose an output-feedback DeePO method to mitigate oscillations in power converter systems. To this end, we first reformulate the output-feedback control problem as a state-feedback LQR problem with a controllable non-minimal state, which can be constructed from input-output signals. This is achieved by the state reduction method in [26]. Then, we apply DeePO to this controllable non-minimal realization to achieve output-feedback adaptive control. Finally, we apply the output-feedback DeePO algorithm to stabilize power converter and direct-drive wind generator systems, both of which are unknown, state-unmeasurable, and encounter sudden change of dynamics due to grid changes. Simulation results show that output-feedback DeePO enables efficient online adaptation and effectively prevents instabilities in grid-connected power converters. Compared with our previous works applying data-enabled predictive control (DeePC) to power converters [21, 27], DeePO offers a significantly reduced online computational burden, making it more suitable for scenarios where the processor cannot solve a quadratic program in real time.
The rest of the paper is organized as follows. Section II formulates the stabilization problem of power converter systems. Section III proposes a controllable non-minimal realization of the output feedback system. Section IV proposes an output-feedback DeePO algorithm for adaptive control. Section V performs simulations on the converter systems. Conclusions are made in Section VI.
@@ II-A Power converter systems
where $x_{t}\in\mathbb{R}^{n}$ is the state variable (could be unmeasurable, e.g., variables in the power grid side), $u_{t}\in\mathbb{R}^{m}$ is the control input (e.g., additional signals added to the current references; see Fig. 1), $y_{t}\in\mathbb{R}^{p}$ is the output (e.g., active and reactive power in our setting), $d_{t}$ is the process noise, and $v_{t}$ is the measurement noise. The unknown system $(A,B,C)$ is controllable and observable, but it may be subject to sudden impedance changes due to external grid changes, e.g., changes of $Z_{\rm grid}$ in Fig. 1. Such events may lead to poorly damped or even unstable oscillations.
@@ II-B Challenges in stabilization of power converter systems (DC source + DC-AC converter + AC power grid)
  • unknown model: Power converter systems often exhibit high-order characteristics, rendering it difficult to establish exact dynamical models. Moreover, the power grid is also unknown when designing a controller for the converter, as the grid is large-scale and time-varying.
@@ II-C Our approach
In this paper, we propose a direct data-driven method to solve the stabilization problem of power converter systems without SysID. Our approach is based on data-enabled policy optimization (DeePO), an adaptive linear quadratic control method introduced in our recent work [23, 24]. It is data-driven and does not involve any explicit SysID. Moreover, DeePO uses online closed-loop data to adaptively update the control policy, making it suitable to deal with time-varying dynamics.
@@
However, the DeePO method [24] works only on the state-feedback setting, which is not the case for the system (1). Leveraging the results from [26], we first propose a controllable non-minimal realization of (1), whose state is measurable from past input-output signals. Then, we propose an output-feedback DeePO algorithm for this non-minimal realization. Finally, we discuss its implementation on power converter systems and perform simulations to validate the effectiveness of the output-feedback DeePO algorithm.
III A controllable non-minimal realization of the output-feedback system
In this section, we leverage the results from [26] to find a controllable non-minimal realization of (1), where the state can be measured from input-output signals.
III-A A non-minimal controllable state
We first consider the system (1) without noise, i.e., $d_{t}=0,v_{t}=0$. We assume that the system order $n$ and lag $l$ is unknown, but we have prior knowledge on an upper bound of the lag $\bar{l}\geq l$. Since the state is unmeasurable, we represent (1) with a non-minimal realization using input and output signals. Denote the input trajectory, the output trajectory, and their stack from time $t-\bar{l}$ to $t-1$ as
${u}_{t,\bar{l}}=\begin{bmatrix}u_{t-1}\\ \vdots\\ u_{t-\bar{l}}\end{bmatrix},~{}{y}_{t,\bar{l}}=\begin{bmatrix}y_{t-1}\\ \vdots\\ y_{t-\bar{l}}\end{bmatrix},~{}{\xi}_{t}=\begin{bmatrix}{u}_{t,\bar{l}}\\ {y}_{t,\bar{l}}\end{bmatrix},$
respectively. Define the extended observability matrix
$\mathcal{O}=\begin{bmatrix}CA^{\bar{l}-1}\\ \vdots\\ CA\\ C\end{bmatrix},$
the controllability matrix ${}\mathcal{C}=[B~{}~{}AB~{}~{}\cdots~{}~{}A^{\bar{l}-1}B]$, and the Toeplitz matrices capturing the impulse response
$\mathcal{T}=\begin{bmatrix}0&CB&CAB&\cdots&CA^{\bar{l}-2}B\\ 0&0&CB&\cdots&CA^{\bar{l}-3}B\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&\cdots&&0&CB\\ 0&0&0&0&0\end{bmatrix}.$
Let $S=[C(\mathcal{C}-A^{\bar{l}}O^{\dagger}\mathcal{T}),CA^{\bar{l}}O^{\dagger}]$. Then, we have the following results.
Lemma 1 (A non-minimal realization)
A non-minimal realization of (1) is given by (LABEL:equ:nonnimimal_realization) shown at the bottom of this page.
Proof:
The state can be represented with system dynamics and past input-output trajectories as
$x_{t}$ $=A^{\bar{l}}x_{t-{\bar{l}}}+\mathcal{C}{u}_{t,{\bar{l}}}$ (2)
${y}_{t,{\bar{l}}}$ $=\mathcal{O}x_{t-{\bar{l}}}+\mathcal{T}{u}_{t,{\bar{l}}}.$
Since ${\bar{l}}\geq l$, the extended observability matrix $\mathcal{O}$ has full column rank, and it has a unique left pseudo inverse $\mathcal{O}^{\dagger}=(\mathcal{O}^{\top}\mathcal{O})^{-1}\mathcal{O}^{\top}$. Then, it follows immediately from (2) that $x_{t}=(\mathcal{C}-A^{\bar{l}}\mathcal{O}^{\dagger}\mathcal{T}){u}_{t,{\bar{l}}}+A^{\bar{l}}\mathcal{O}^{\dagger}{y}_{t,{\bar{l}}}$ and $y_{t}=Cx_{t}=S\xi_{t}.$ Thus, a non-minimal realization of (1) is given by (LABEL:equ:nonnimimal_realization).
However, for multiple-output systems with $p>1$, the non-minimal realization (LABEL:equ:nonnimimal_realization) is generally not controllable. Moreover, the corresponding input-state data matrix
$\begin{bmatrix}u_{0}&u_{1}&\dots&u_{t-1}\\ \xi_{0}&\xi_{1}&\dots&\xi_{t-1}\end{bmatrix}\in\mathbb{R}^{(m(\bar{l}+1)+p\bar{l})\times t}$ (3)
will never have full row rank [26], which precludes the application of DeePO on the non-minimal realization (LABEL:equ:nonnimimal_realization) [24]. In fact, it can be shown that the maximal rank of (3) can only be $m(\bar{l}+1)+n$ [26].
To ensure the full rank condition of data, we adopt the approach [26] that constructs an reduced non-minimal state for (1). The following lemma is a direct implication of [26, Theorem 4].
Lemma 2 (A controllable non-minimal realization)
Let $d_{t}=0,v_{t}=0$ for (1). Then, there exists a full row rank permutation matrix $T\in\mathbb{R}^{(m\bar{l}+n)\times(m\bar{l}+p\bar{l})}$ and $(A_{z},B_{z},C_{z})$ such that a controllable non-minimal state-space representation is given by
$z_{t+1}$ $=A_{z}z_{t}+B_{z}u_{t}$ (4)
$y_{t}$ $=C_{z}z_{t}$
with a non-minimal state
$z_{t}=T\xi_{t}.$ (5)
Since the realization (4) is controllable, we can choose persistently exciting (PE) inputs of order $\bar{l}+1+n$ such that
$D_{0,t}=\begin{bmatrix}u_{0}&u_{1}&\dots&u_{t-1}\\ z_{0}&z_{1}&\dots&z_{t-1}\end{bmatrix}\in\mathbb{R}^{(m(\bar{l}+1)+n)\times t}$ (6)
has full row rank [26]. Hence, we can design data-driven state-feedback controller for (4), where $z_{t}$ can be computed with the permutation $T$ and input-output signals.
Next, we show how to obtain the permutation matrix $T$ from a batch of input-output data [26].
III-B Computing the permutation matrix from input-output data
When there is no noise in the system (1), the permutation matrix $T$ can be solved directly from a batch of input-output data [26]. Specifically, suppose that we have the data generated by PE input of order $\bar{l}+1+n$
$\Xi_{0,t}=\begin{bmatrix}{\xi}_{0}&{\xi}_{1}&\dots&{\xi}_{t-1}\end{bmatrix}\in\mathbb{R}^{(m\bar{l}+p\bar{l})\times t}.$ (7)
Then, $T\in\mathbb{R}^{(m\bar{l}+n)\times(m\bar{l}+p\bar{l})}$ is the permutation matrix such that $T\Xi_{0,t}$ has full row rank $m\bar{l}+n$. When $n$ is unknown, we can easily separate linearly independent rows by, e.g., Gaussian elimination, and further obtain $T$. Moreover, the system order $n$ can be obtained as the number of rows in $T$.
When the noise $d_{t}$ and $v_{t}$ in system (1) are not zero, the input-state data matrix (3) of system (LABEL:equ:nonnimimal_realization) can be full row rank. In this case, we use singular value decomposition (SVD) for $\Xi_{0,t}$
$\Xi_{0,t}=\begin{bmatrix}U_{r}&U_{l}\end{bmatrix}\begin{bmatrix}\Lambda_{r}&0\\ 0&\Lambda_{l}\end{bmatrix}\begin{bmatrix}V_{r}^{\top}\\ V_{l}^{\top}\end{bmatrix},$
where $\Lambda_{r}$ is the singular value matrix with largest $r$ singular values. The number of rows $r$ of the permutation matrix $T$ should be $m\bar{l}+n$, which is consistent with Lemma 2. Since $n$ is unknown, we can select $r>m\bar{l}$ such that there is a clear distinction between the largest $r$ singular values and the remaining ones (corresponding to noise). Then, the permutation matrix $T$ is given by
$T=\Lambda_{r}^{-1}U_{r}^{\top},$
i.e., it selects the orthogonal basis for the row space of $\Xi_{0,t}$ corresponding to largest $r$ singular values, i.e., $V_{r}^{\top}=T\Xi_{0,t}.$
Once we obtain the permutation matrix $T$, we can measure the non-minimal state $z_{t}$ from input-output signals via (5). Moreover, the input-state data matrix (6) has full rank, which enables us to use DeePO for the realization (4) for adaptive control.
@@ III-B Computing the permutation matrix from input-output data
IV Output-feedback data-enabled policy optimization for direct adaptive control
In this section, we first introduce the data-driven LQR formulation with covariance parameterization for the non-minimal realization (4). Based on this, we propose a DeePO method for direct adaptive control of (1) using input-output data.
IV-A Direct data-driven LQR design of (4) with covariance parameterization
Consider the controllable non-minimal realization (4) with noise
@@ IV-A Direct data-driven LQR design of (4) with covariance parameterization
Here, $h_{t}$ is the performance signal of interest and the weighting matrices $(Q,R)$ are positive definite.
The LQR problem is phrased as finding a state-feedback gain $K\in\mathbb{R}^{m\times(m\bar{l}+n)}$ that minimizes the $\mathcal{H}_{2}$-norm of the transfer function $\mathscr{T}(K):w\rightarrow h$ of the closed-loop system
$\begin{bmatrix}z_{t+1}\\ h_{t}\end{bmatrix}=\begin{bmatrix}A_{z}+B_{z}K&I_{n}\\ \hline\cr\begin{bmatrix}Q^{1/2}\\ R^{1/2}K\end{bmatrix}&0\end{bmatrix}\begin{bmatrix}z_{t}\\ d_{t}\end{bmatrix}.$
When $A_{z}+B_{z}K$ is stable, it holds that [28]
$\|\mathscr{T}(K)\|_{2}^{2}=\text{Tr}((Q+K^{\top}RK)\Sigma_{K})=:J(K),$ (9)
@@
$\Sigma_{K}=I_{n}+(A_{z}+B_{z}K)\Sigma_{K}(A_{z}+B_{z}K)^{\top}.$ (10)
We refer to $J(K)$ as the LQR cost and to (9)-(10) as a policy parameterization of the LQR. When the model parameters are known, the optimal LQR gain $K^{*}$ is unique and can be found by, e.g., solving an algebraic Riccati equation [28].
Since $(A_{z},B_{z})$ is unknown, data-driven methods learn the LQR gain from input-state data. Consider the $t$-long time series of states, inputs, noises, and successor states of (4)
$Z_{0,t}$ $:=\begin{bmatrix}z_{0}&z_{1}&\dots&z_{t-1}\end{bmatrix}\in\mathbb{R}^{(m\bar{l}+n)\times t},$
@@
$W_{0,t}$ $:=\begin{bmatrix}w_{0}&w_{1}&\dots&w_{t-1}\end{bmatrix}\in\mathbb{R}^{(m\bar{l}+n)\times t},$
$Z_{1,t}$ $:=\begin{bmatrix}z_{1}&z_{2}&\dots&z_{t}\end{bmatrix}\in\mathbb{R}^{(m\bar{l}+n)\times t},$
@@
$Z_{1,t}=A_{z}Z_{0,t}+B_{z}U_{0,t}+W_{0,t}.$ (11)
Assume that the input data is PE of order $\bar{l}+1+n$. Then, the block matrix $[U_{0,t}^{\top},Z_{0,t}^{\top}]^{\top}$ has full row rank. Define the sample covariance of input-state data as
$\Phi_{t}:=\frac{1}{t}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}^{\top},$ (12)
which is positive definite due to the full rank condition. Then, we can use sample covariance to parameterize the policy
$\begin{bmatrix}K\\ I_{(m\bar{l}+n)}\end{bmatrix}=\Phi_{t}V,$ (13)
where $V\in\mathbb{R}^{(m\bar{l}+m+n)\times(m\bar{l}+m+n)}$.
With the covariance parameterization (13), the LQR problem (9)-(10) can be expressed by raw data matrices $(Z_{0,t},U_{0,t},Z_{1,t})$ and the optimization matrix $V$. For brevity, let $\overline{Z}_{0,t}=Z_{0,t}D_{0,t}^{\top}/t$ and $\overline{U}_{0,t}=U_{0,t}D_{0,t}^{\top}/t$ be a partition of $\Phi_{t}$, and let $\overline{W}_{0,t}=W_{0,t}D_{0,t}^{\top}/t$ be the noise-state-input covariance, and finally define the covariance with respect to the successor state $\overline{Z}_{1,t}=Z_{1,t}D_{0,t}^{\top}/t$. Then, the closed-loop matrix can be written as
@@
Following the certainty-equivalence principle [14], we disregard the unmeasurable $\overline{W}_{0,t}$ and use $\overline{Z}_{1,t}V$ as the closed-loop matrix. After substituting $A+BK$ with $\overline{Z}_{1,t}V$ in (9)-(10) and leveraging (13), the LQR problem becomes
$\mathop{\text{minimize}}\limits_{V}~{}J_{t}(V):=\text{Tr}\left((Q+V^{\top}\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}V)\Sigma_{t}\right),$ (14)
subject $\text{subject to}~{}~{}\overline{Z}_{0,t}V=I_{m\bar{l}+n},$
where $\Sigma_{t}=I_{m\bar{l}+n}+\overline{Z}_{1,t}V\Sigma_{t}V^{\top}\overline{Z}_{1,t}^{\top}$ is a covariance parameterization of (10), and the gain matrix can be recovered by $K=\overline{U}_{0,t}V$. We refer to (14) as the direct data-driven LQR problem, which does not involve any explicit SysID.
IV-B Output-feedback DeePO
Policy optimization (PO) refers to a class of direct design methods, where the policy is parameterized and recursively updated using gradient methods [29]. In particular, the DeePO algorithm uses online gradient descent of (14) to recursively update $V$. The details are presented in Algorithm 1. Given offline data, we first compute the permutation matrix $T$. At time $t$, we apply the linear state feedback policy $u_{t}=K_{t}z_{t}+e_{t}$ for control, where $e_{t}\in\mathbb{R}^{m}$ is a probing noise used to ensure the PE rank condition. We use online projected gradient descent to update the parameterized policy, where the projection $\Pi_{\overline{Z}_{0,t+1}}:=I_{m(\bar{l}+1)+n}-\overline{Z}_{0,t+1}^{\dagger}\overline{Z}_{0,t+1}$ onto the nullspace of $\overline{Z}_{0,t+1}$ is to ensure the subspace constraint in (14). Define the feasible set of (14) (i.e., the set with stable closed-loop matrices) as $\mathcal{S}_{t}:=\{V\mid\overline{Z}_{0,t}V=I_{m\bar{l}+n},\rho(\overline{Z}_{1,t}V)<1\}$. The gradient can be computed by the following lemma.
Lemma 3 ([24, Lemma 2])